40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 24 15:24:13 2021, gautam, Update, LSC, Notes on tests 
    Reply  Wed Mar 24 22:54:49 2021, gautam, Update, LSC, New day, new problems PRMI3f_noArmssensMat.pdfREFL55_whtGainStepping.pngREFL55_whtGainStepping2.png
       Reply  Mon Mar 29 19:32:46 2021, gautam, Update, LSC, REFL55 whitening checkout REFL55wht.png
Message ID: 15977     Entry time: Mon Mar 29 19:32:46 2021     In reply to: 15960
Author: gautam 
Type: Update 
Category: LSC 
Subject: REFL55 whitening checkout 

I repeated the usual whitening board characterization test of:

  • driving a signal (using awggui) into the two inputs of the whitening board using a spare DAC channel available in 1Y2
  • demodulating the response using the LSC sensing matrix infrastructure
  • Stepping the whitening gain, incrementing it in 3dB steps, and checking if the demodulated lock-in outputs increase in the expected 3dB steps.

Attachment #1 suggests that the steps are equal (3dB) in size, but note that the "Q" channel shows only ~half the response of the I channel. The drive is derived from a channel of an unused AI+dewhite board in 1Y2, split with a BNC Tee, and fed to the two inputs on the whitening filter. The impedance is expected to be the same on each channel, and so each channel should see the same signal, but I see a large asymmetry. All of this checked out a couple of weeks ago (since we saw ellipses and not circles) so not sure what changed in the meantime, or if this is symptomatic of some deeper problem.

Usually, doing this and then restoring the cabling returns the signal levels of REFL55 to nominal levels. Today it did not - at the nominal whitening gain setting of +18dB flat gain, when the PRMI is fringing, the REFL55 inputs are frequently reporting ADC overflows. Needless to say, all my attempts today evening to transition the length control of the vertex from REFL165 to REFL55 failed.

I suppose we could try shifting the channels to (physical) Ch5 and Ch6 which were formerly used to digitize the ALS DFD outputs and are currently unused (from Ch3, Ch4) on this whitening filter and see if that improves the situation, but this will require a recompile of the RTCDS model and consequent CDS bootfest, which I'm not willing to undertake today. If anyone decides to do this test, let's also take the opportunity to debug the BIO switching for the delay line.

Attachment 1: REFL55wht.png  22 kB  Uploaded Mon Mar 29 20:45:25 2021  | Hide | Hide all
REFL55wht.png
ELOG V3.1.3-